OpenStudy (anonymous):

I need help with: The sum of the first 10 term of an Arithmetic.Progression is 120 and the sum of the first twenty terms is 840. Find the sum of the first thirty terms.

OpenStudy (anonymous):

10(2a+9r)/2=120 20(2a+19r)/2=840 solve that and use a and r

OpenStudy (anonymous):

quetejedi, shouldn't it be 5 and 10, the formula is N/2?

OpenStudy (anonymous):

there is it, i divided by 2

OpenStudy (anonymous):

\[\S_{n}=n/2(2a+(n-1)d)\]120=5(2a+9d) 840=10(2a+19d) 120=10a+45d 840=20a+15960d 240=20a+90d 840=20a+15960d d=26.45 a=-107.025 gues this can help

OpenStudy (anonymous):

Thank you very much, this helped me. I think you have 840=20a+15960d wrong it should be 190 because of 840=10(2a+19d)

OpenStudy (anonymous):

I did it like this: S10 = 5 ( 2a + 9d) = 120 S20 = 10 (2a + 19d) = 840 2a +9d = 120/5 = 24 2a + 19d = 840 / 10 = 84 then 10d (from 19-9d) = 84 + 24 = 60 d=6 a = -15 530 = 15 (2* - 15 + 29 * 6 = 2160?

OpenStudy (anonymous):

yap